I am lost on this math. Please help. Thank you
f(x)=-2x^2+2x+3
The x-coordinate of the vertex is.....

Answer by edjones(8007)

f(x)=-2x^2+2x+3
x coordinate of the vertex is -b/2a
a=-2, b=2
-b/2a=-2/2*-2
=-2/-4
=1/2 Ans.
.
Ed